OpenStudy (lukecrayonz):

Find the equation of the tangent line of the graph f(x)=1/x^5 at: (1,1)

OpenStudy (lukecrayonz):

Derivative is -5/x^6

OpenStudy (anonymous):

Take the first derivative to fine the slope at that point. once you have the slope you can find the y-intercept by using y=mx+b where m is the slope that you obtained from the second derivative.

OpenStudy (anonymous):

I'm sorry first derivative which is the one you provided there.

OpenStudy (lukecrayonz):

@perdomo.arturo so now I input (1,1) for y and x and solve for the slope? or.. :O

OpenStudy (anonymous):

yeah use that point to find the slope. dy/dx is equal to m or the slope.

OpenStudy (anonymous):

Really you just need the x value for the slope, but you'll need both to find the y-intercept of the line using the equation I provided there.

OpenStudy (lukecrayonz):

So then 1=-5/1^6?

OpenStudy (lukecrayonz):

No no wait lol

OpenStudy (lukecrayonz):

y=-5/x^6+b b=6

OpenStudy (anonymous):

almost right. y=-5x+6. Remeber that a tangent line is just a straight line.